# Gradient Existence and Energy Finiteness of Local Minimizers in the Wasserstein $L^{\infty}$ Topology for Binary-Star Systems

### v1

@@ 4 Wasserstein $L^{\infty}$ Metric on ${R}\left(\mathbb{R}^{3}\right)$
Notice when $R\rightarrow\infty$, $0\leq\mu\left(\{x\in E\mid\rho(x)>R\}\leq\mu(\{x\in\mathbb{R}^{3}\mid\rho(x)>R\right\}$ $\leq\sum_{N=[R]}^{\infty}\mu(\{x\mid N\leq\rho(x)<N+1\})\rightarrow 0$, hence $\mu(\{x\in E\mid\rho(x)\leq R\})=\mu(E)-\mu(\{x\in E\mid\rho(x)>R\}\rightarrow\mu(E)$. By similar arguments in the following paragraph about how to get $\widetilde{E}$ from $\widetilde{E}_{k}$, we can show when $R$ is large enough, $\mu(\{x\in E\mid\rho(x)\leq R\})>\frac{1}{2}\mu(E)>0$.
Pick one of such $R$, and let $\widetilde{E}_{k}=\{x\in E\mid\rho(x)\leq R,|x|\leq k\}$, then $\left\{\widetilde{E}_{k}\right\}_{k=1}^{\infty}$ is an ascending collection of measurable sets, and $\cup_{k=1}^{\infty}\widetilde{E}_{k}=\{x\in E\mid\rho(x)\leq R\}$, therefore, $\mu(\{x\in E\mid\rho(x)\leq R\})=\lim\limits_{k\rightarrow\infty}\mu\left(\widetilde{E}_{k}\right)$. (see [34, Section 2.5]). When $k$ is large enough, $\mu(\{x\in E\mid\rho(x)\leq R,|x|\leq k\})>\frac{1}{2}\mu(E)>0$, then replace $R$ by another larger $R$ if needed, we have $\mu(\{x\in E\mid\rho(x)\leq R,|x|\leq R\})>0$. We set $\widetilde{E}=\{x\in E\mid\rho(x)\leq R,|x|\leq R\}$, then $\sigma(x)=\left\{\begin{array}[]{ll}1,&x\in\widetilde{E}\\ 0,&\text{otherwise }\end{array}\right$. has the properties we say above. In particular, one can see $0\in P_{\infty}(\rho)$, but $P_{\infty}(\rho)\neq\{0\}$. Notice in $\widetilde{E}$, $\rho(x)$ is bounded and $0\leq\rho(x)\leq R$.
@@ 5 Analysis of Lagrange Multiplier and Properties of Local Minimizers
This section is divided into three subsections. In the first subsection, we compute the variational derivative of $E_{J}$, and prove the finite energy property of local minimizers, which ensures variational derivative makes sense. In the second subsection, we revisit and modify the proofs of “weak” versions of Theorem 2.25 (iviv-$v$). Here “weak” means (EL) holds true for almost all points, rather than pointwise, in the region we consider. In the third subsection, we point out the particularity of the topology induced by $W^{\infty}$ metric by discussing minimizers’ existence and nonexistence.
@@ 5.1 Variational derivative of $E_{J}$ and finite energy of local minimizer
In order to consider the variational derivative $E_{J}^{\prime}(\rho)$ of the energy $E_{J}(\rho)$, we first clarify the space where we want to take derivatives: $X=L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\cap L^{1}\left(\mathbb{R}^{3}\right)$, $U=\{\rho\in X\mid\rho\geq 0,U(\rho)<\left.\infty,G(\rho,\rho)<\infty,T_{J}(\rho)<\infty\right\}$, $W=\left\{\rho\in U\mid\int\rho=1\right\}$.
@@
Suppose $\rho\in X$ is non-negative and positive mass, then by Remark 2.9 we know $I(\rho)>0$, thus $T_{J}(\rho)$ is finite. Thanks to Proposition 2.10 we have $G(\rho,\rho)<\infty$. Therefore,
@@
$U(\sigma)=\int A(\sigma(x))d^{3}x=\int_{\{\sigma<\delta\}}A(\sigma(x))d^{3}x+\int_{\{\delta\leq\sigma\leq C\}}A(\sigma(x))d^{3}x=:U_{1}+U_{2}$
where $C=\|\sigma\|_{L^{\infty}(\mathbb{R}^{3})}$. By the choice of $\delta$ we know $U_{1}\leq\int_{\{\sigma<\delta\}}\sigma^{\frac{4}{3}}(x)d^{3}x\leq\int\sigma^{\frac{4}{3}}(x)d^{3}x$, Since $\sigma\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{\infty}\left(\mathbb{R}^{3}\right)$, we know by Interpolation Inequality Proposition A.6 that $U_{1}\leq\int\sigma^{\frac{4}{3}}(x)d^{3}x<\infty$. Since $A(s)$ is increasing, we know
$U_{2}$ $\leq\int_{\{\delta\leq\sigma\leq C\}}A(C)d^{3}x$
$\leq A(C)\int_{\{\delta\leq\sigma\leq C\}}\frac{\sigma(x)}{\delta}d^{3}x$
$\leq\frac{A(C)}{\delta}\int_{\{\delta\leq\sigma\leq C\}}{\sigma(x)}d^{3}x$
@@
Given $\rho\in W$, we have $P_{\infty}(\rho)\subset P(\rho)$, and $E_{J}(\rho)$ is $P_{\infty}(\rho)$-differentiable at $\rho$. Here $P(\rho)$ and $P_{\infty}(\rho)$-differentiability (i.e. differentiability at $\rho$ in the direction of $P_{\infty}(\rho)$) are provided in Appendix A.1. Moreover, the derivative at $\rho$ is $E_{J}^{\prime}(\rho)$ in the sense that $\forall\sigma\in P_{\infty}(\rho)$, $E_{J}^{\prime}(\rho)(\sigma)=\int E_{J}^{\prime}(\rho)\sigma$ [^3] To remain consistent with the notation in [5, Section 4], we use $E_{J}^{\prime}$ as the symbol for both linear functional and function, provided it does not cause confusion., where on the right hand side the function $E_{J}^{\prime}(\rho)$ is given by
@@
Given $\sigma\in P_{\infty}(\rho)$, then $\exists R>0$ such that $\sigma\in P_{R}(\rho)$. By construction we know $\sigma\in L^{\infty}\left(\mathbb{R}^{3}\right)$ with compact support, thus by Hölder’s inequality we know $\sigma\in L^{p}\left(\mathbb{R}^{3}\right)$ for $p>0$. Since $\rho\in W$, $\int\rho=1$, then when $t$ is small enough, we have $\int\rho+t\sigma>0$. Thanks to Remark 4.12 and Remark 5.1, we know $\rho+t\sigma$ is nonnegative, $T_{J}(\rho+t\sigma)<\infty$ and $G(\rho+t\sigma,\rho+t\sigma)<\infty$. Moreover, to show $\rho+t\sigma\in U$, it suffices to show $U(\rho+t\sigma)<\infty$.
@@
The right hand side is an integrable function. Then we know $U(\rho+t\sigma)=\int A(\rho+t\sigma)<\infty$. Then we know $\rho+t\sigma\in U$ for $t$ small enough. Therefore, $\sigma\in P(\rho)$ and then $P_{\infty}(\rho)\subset P(\rho)$.
The observations above are also helpful to calculate the derivative. We notice $|\frac{A(\rho(x)+t\sigma(x))-A(\rho(x))}{t}|$ is also bounded by an integrable function $A^{\prime}(R+\|\sigma\|_{L^{\infty}})\|\sigma\|_{L^{\infty}}\cdot\mathbf{1}_{E_{R}}$. Moreover, by (29) we know $\frac{A(\rho(x)+t\sigma(x))-A(\rho(x))}{t}$ converges a.e. to $A^{\prime}(\rho(x))\sigma(x)$ as $t\rightarrow 0$. By dominated convergence theorem, we have
$\lim_{t\rightarrow\infty}\frac{\int A(\rho+t\sigma)-\int A(\rho)}{t}=\int A^{\prime}(\rho)\cdot\sigma.$ (30)
Now consider gravitational interaction energy. Due to Remark 2.11, we know $G(\rho,\rho)=\int\rho V_{\rho}<\infty$ and $G(\rho+t\sigma,\rho+t\sigma)<\infty$, then we compute
@@
Since $G(,\cdot,)$ is a symmetric bilinear form, thus $G(\rho,\sigma)+G(\sigma,\rho)=2G(\sigma,\rho)=2\int\sigma V_{\rho}$.
Consider $T_{J}(\rho)=\frac{J^{2}}{2I(\rho)}$, when $t$ is small enough, we have $\int\rho+t\sigma>0$, thus $I(\rho+t\sigma)>0$. and $T_{J}(\rho+t\sigma)$ is well-defined. Easy to check the mass of $t\sigma$ is $m(t\sigma)=tm(\sigma)$, the center of mass is $\bar{x}(t\sigma)=\bar{x}(\sigma)$, and the moment of inertia is $I(t\sigma)=tI(\sigma)$. Thanks to Lemma 2.15, we can expand $I(\rho+t\sigma)$ and get
@@
$I(\sigma)+m(\sigma)\cdot r^{2}(\bar{x}(\rho)-\bar{x}(\sigma))=$ $\int\sigma\cdot\left(r^{2}(x-\bar{x}(\sigma))+r^{2}(\bar{x}(\rho)-\bar{x}(\sigma))\right)$
$=$ $\int\sigma\cdot(\langle x,x\rangle_{2}-2\langle\bar{x}(\sigma),x\rangle_{2}+\langle\bar{x}(\sigma),\bar{x}(\sigma)\rangle_{2}$
@@
$-2\langle\bar{x}(\rho),x\rangle_{2}+2\langle\bar{x}(\rho),x\rangle_{2})$
$=$ $\int\sigma\cdot(r^{2}(x-\bar{x}(\rho))+2\langle\bar{x}(\sigma),\bar{x}(\sigma)\rangle_{2}$
$-2\langle\bar{x}(\sigma),x\rangle_{2}-2\langle\bar{x}(\sigma),\bar{x}(\rho)\rangle_{2}+2\langle\bar{x}(\rho),x\rangle_{2})$
$=$ $\int\sigma\cdot r^{2}(x-\bar{x}(\rho))+2m(\sigma)\langle\bar{x}(\sigma),\bar{x}(\sigma)\rangle_{2}-2m(\sigma)\langle\bar{x}(\sigma),\bar{x}(\sigma)\rangle_{2}$
@@
$=$ $\int r^{2}(x-\bar{x}(\rho))\cdot\sigma$
Therefore, $\frac{T_{J}(\rho+t\sigma)-T_{J}(\rho)}{t}\rightarrow\int\frac{-J^{2}}{2I^{2}(\rho)}\cdot r^{2}(x-\bar{x}(\rho))\cdot\sigma$ as $t\rightarrow 0$. Collect the results above we finish the proof of Lemma 5.3. ∎
@@
The admissible perturbations are a bit different from $P_{0}(\rho)$ mentioned in [5, Section 4]. There the energy related to momentum is not $T_{J}(\rho)$ but $\frac{1}{2}\int\rho(x)L\left(m_{\rho}(r(x))\right)r^{-2}(x)d^{3}x$, in which the term $r^{-2}(x)$ will be singular when $r(x)\rightarrow 0$. Therefore, the authors require $\sigma(x)=0$ when $r(x)$ is small.
Remark 5.6.
Notice that we rely on the fact that $\sigma$ belongs to $L^{\infty}$ to apply the Dominated Convergence Theorem to obtain (30). Without this assumption, one might consider using Fatou’s Lemma instead, but that would only provide boundedness rather than convergence.
Since an element $\rho\in R(\mathbb{R}^{3})$ might have infinite total energy $E_{J}(\rho)$ — equivalently, infinite internal energy $U(\rho)$ by Remark 5.1 — we have $W\subsetneq R(\mathbb{R}^{3})$. A simple example with infinite total energy could be $\rho(x)\coloneq\rho_{N}(x)=C\frac{1}{|x|^{\alpha}}\mathbf{1}_{\{{{|x|}<N}\}}$ with suitable $C$, $\alpha$, and $N$. Nevertheless, we can show if $\rho$ is a local minimizer, then $\rho\in W$.
@@
If $\rho$ minimizes $E_{J}(\rho)$ locally on ${R}\left(\mathbb{R}^{3}\right)$, then $\rho\in W$. In particular, by Lemma 5.3, we know $E_{J}(\rho)$ is $P_{\infty}(\rho)$-differentiable at $\rho$, and the derivative at $\rho$ is $E_{J}^{\prime}(\rho)$ given in (28).
@@
Since $\int\rho=1>0$, similar to Remark 5.1 we know $T_{J}(\rho)$ and $G(\rho,\rho)$ are finite. Since $\rho$ is a local minimizer, $\exists\delta>0$, such that $\forall\sigma\in{R}\left(\mathbb{R}^{3}\right)$ with $W^{\infty}(\rho,\sigma)<\delta$, one has $E_{J}(\rho)\leq E_{J}(\sigma)$. Thanks to Lemma 4.7 (vivi), take $\delta$ small enough, we can find a $\sigma\in{R}\left(\mathbb{R}^{3}\right)\cap L^{\infty}\left(\mathbb{R}^{3}\right)$, with $W^{\infty}(\rho,\sigma)<\delta$ and $\int\sigma>0$. Similar as above we have $G(\sigma,\sigma)<\infty,T_{J}(\sigma)<\infty$. Thanks to Lemma 5.2, we know $U(\sigma)<\infty$, and then $E_{J}(\rho)\leq E_{J}(\sigma)<\infty$, which implies $U(\rho)<\infty$. Then we have $\rho\in W$. ∎

### v2

@@ 4 Wasserstein $L^{\infty}$ Metric on ${R}\left(\mathbb{R}^{3}\right)$
Notice when $R\rightarrow\infty$,
$0\leq\mu\left(\{x\in E\mid\rho(x)>R\}\leq\mu(\{x\in\mathbb{R}^{3}\mid\rho(x)>R\right\}\leq\sum_{N=[R]}^{\infty}\mu(\{x\mid N\leq\rho(x)<N+1\})\rightarrow 0$
Hence
$\mu(\{x\in E\mid\rho(x)\leq R\})=\mu(E)-\mu(\{x\in E\mid\rho(x)>R\}\rightarrow\mu(E)$
By similar arguments in the following paragraph about how to get $\widetilde{E}$ from $\widetilde{E}_{k}$, we can show when $R$ is large enough, $\mu(\{x\in E\mid\rho(x)\leq R\})>\frac{1}{2}\mu(E)>0$.
Pick one of such $R$, and let $\widetilde{E}_{k}=\{x\in E\mid\rho(x)\leq R,|x|\leq k\}$, then $\left\{\widetilde{E}_{k}\right\}_{k=1}^{\infty}$ is an ascending collection of measurable sets, and $\cup_{k=1}^{\infty}\widetilde{E}_{k}=\{x\in E\mid\rho(x)\leq R\}$, therefore, $\mu(\{x\in E\mid\rho(x)\leq R\})=\lim\limits_{k\rightarrow\infty}\mu\left(\widetilde{E}_{k}\right)$. (see [34, Section 2.5]). When $k$ is large enough,
$\mu(\{x\in E\mid\rho(x)\leq R,|x|\leq k\})>\frac{1}{2}\mu(E)>0$
Then replace $R$ by another larger $R$ if needed, we have $\mu(\{x\in E\mid\rho(x)\leq R,|x|\leq R\})>0$. We set $\widetilde{E}=\{x\in E\mid\rho(x)\leq R,|x|\leq R\}$, then $\sigma(x)=\left\{\begin{array}[]{ll}1,&x\in\widetilde{E}\\ 0,&\text{otherwise }\end{array}\right$. has the properties mentioned at the beginning of this remark. In particular, one can see $0\in P_{\infty}(\rho)$, but $P_{\infty}(\rho)\neq\{0\}$. Notice in $\widetilde{E}$, $\rho(x)$ is bounded and $0\leq\rho(x)\leq R$.
@@ 5 Analysis of Lagrange Multiplier and Properties of Local Minimizers
This section is divided into three subsections. In the first subsection, we compute the variational derivative of $E_{J}$, and prove the finite energy property of local minimizers, which ensures the variational derivative makes sense. In the second subsection, we revisit and modify the proofs of “weak” versions of Theorem 2.22 (iviv-$v$). Here “weak” means (EL) holds true for almost all points, rather than pointwise, in the region we consider. In the third subsection, we discuss the nonexistence of finite energy local minimizers under the topology inherited from a topological vector space, thereby suggesting the particularity of the topology induced by the $W^{\infty}$ metric.
@@ 5.1 Variational derivative of $E_{J}$ and finite energy of local minimizer
Note that an element $\rho\in R(\mathbb{R}^{3})$ might have infinite total energy $E_{J}(\rho)$. A simple example with infinite total energy could be $\rho(x)\coloneq\rho_{N}(x)=C\frac{1}{|x|^{\alpha}}\mathbf{1}_{\{{{|x|}<N}\}}$ with suitable $C$, $\alpha$, and $N$. In order to consider the variational derivative $E_{J}^{\prime}(\rho)$ of the energy $E_{J}(\rho)$, we first determine the space $W\subsetneq R(\mathbb{R}^{3})$ where we want to take derivatives:
$X=L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\cap L^{1}\left(\mathbb{R}^{3}\right)$
$U=\{\rho\in X\mid\rho\geq 0\text{ , }U(\rho)<\left$$\infty\right\}$
$W=\left\{\rho\in U\mid\int_{\mathbb{R}^{3}}\rho\,dx=1\right\}$
@@
Suppose $\rho\in X$ is non-negative and positive mass, then by Remark 2.8 we know $I(\rho)>0$, thus $T_{J}(\rho)$ is finite. Thanks to Proposition 2.9, we have $G(\rho,\rho)<\infty$. Therefore,
@@
$U(\sigma)=\int_{\mathbb{R}^{3}}A(\sigma(x))\,dx=\int_{\{\sigma<\delta\}}A(\sigma(x))\,dx+\int_{\{\delta\leq\sigma\leq C\}}A(\sigma(x))\,dx=:U_{1}+U_{2}$
where $C=\|\sigma\|_{L^{\infty}(\mathbb{R}^{3})}$. By the choice of $\delta$ we know
$U_{1}\leq\int_{\{\sigma<\delta\}}\sigma^{\frac{4}{3}}(x)\,dx\leq\int_{\mathbb{R}^{3}}\sigma^{\frac{4}{3}}(x)\,dx$
Since $\sigma\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{\infty}\left(\mathbb{R}^{3}\right)$, we know by Interpolation Inequality [9, Section 4.2] that $U_{1}\leq\int_{\mathbb{R}^{3}}\sigma^{\frac{4}{3}}(x)\,dx<\infty$. Since $A(s)$ is increasing, we know
$U_{2}$ $\leq\int_{\{\delta\leq\sigma\leq C\}}A(C)\,dx$
$\leq A(C)\int_{\{\delta\leq\sigma\leq C\}}\frac{\sigma(x)}{\delta}\,dx$
$\leq\frac{A(C)}{\delta}\int_{\{\delta\leq\sigma\leq C\}}{\sigma(x)}\,dx$
@@
Given $\rho\in W$, we have $P_{\infty}(\rho)\subset P(\rho)$, and $E_{J}(\rho)$ is $P_{\infty}(\rho)$-differentiable at $\rho$. Here $P(\rho)$ and $P_{\infty}(\rho)$-differentiability (i.e. differentiability at $\rho$ in the direction of $P_{\infty}(\rho)$) are provided in Appendix A. Moreover, the derivative at $\rho$ is $E_{J}^{\prime}(\rho)$ in the sense that $\forall\sigma\in P_{\infty}(\rho)$, $E_{J}^{\prime}(\rho)(\sigma)=\int_{\mathbb{R}^{3}}E_{J}^{\prime}(\rho)\sigma\,dx$ [^3] To remain consistent with the notation in [5, Section 4], we use $E_{J}^{\prime}$ as the symbol for both linear functional and function, provided it does not cause confusion., where on the right hand side the function $E_{J}^{\prime}(\rho)$ is given by
@@
Given $\sigma\in P_{\infty}(\rho)$, then $\exists R>0$ such that $\sigma\in P_{R}(\rho)$. By construction we know $\sigma\in L^{\infty}\left(\mathbb{R}^{3}\right)$ with compact support, thus by Hölder’s inequality we know $\sigma\in L^{p}\left(\mathbb{R}^{3}\right)$ for $p>0$. Since $\rho\in W$, $\int_{\mathbb{R}^{3}}\rho\,dx=1$, then when $t$ is small enough, we have $\int_{\mathbb{R}^{3}}(\rho+t\sigma)\,dx>0$. Thanks to Remark 4.9 and Remark 5.1, we know $\rho+t\sigma$ is nonnegative, $T_{J}(\rho+t\sigma)<\infty$ and $G(\rho+t\sigma,\rho+t\sigma)<\infty$. Moreover, to show $\rho+t\sigma\in U$, it suffices to show $U(\rho+t\sigma)<\infty$.
@@
The right hand side is an integrable function. Then we know $U(\rho+t\sigma)=\int_{\mathbb{R}^{3}}A(\rho+t\sigma)\,dx<\infty$. Then we know $\rho+t\sigma\in U$ for $t$ small enough. Therefore, $\sigma\in P(\rho)$ and then $P_{\infty}(\rho)\subset P(\rho)$.
The observations above are also helpful to calculate the derivative. We notice $|\frac{A(\rho(x)+t\sigma(x))-A(\rho(x))}{t}|$ is also bounded by an integrable function $A^{\prime}(R+\|\sigma\|_{L^{\infty}})\|\sigma\|_{L^{\infty}}\cdot\mathbf{1}_{E_{R}}$. Moreover, by (30) we know $\frac{A(\rho(x)+t\sigma(x))-A(\rho(x))}{t}$ converges a.e. to $A^{\prime}(\rho(x))\sigma(x)$ as $t\rightarrow 0$. By the dominated convergence theorem, we have
$\lim_{t\rightarrow\infty}\frac{\int_{\mathbb{R}^{3}}A(\rho+t\sigma)\,dx-\int_{\mathbb{R}^{3}}A(\rho)\,dx}{t}=\int_{\mathbb{R}^{3}}A^{\prime}(\rho)\cdot\sigma\,dx$ (31)
Now consider gravitational interaction energy. Due to Remark 2.10, we know $G(\rho,\rho)=\int_{\mathbb{R}^{3}}\rho V_{\rho}\,dx<\infty$ and $G(\rho+t\sigma,\rho+t\sigma)<\infty$, then we compute
@@
Since $G(,\cdot,)$ is a symmetric bilinear form, thus
$G(\rho,\sigma)+G(\sigma,\rho)=2G(\sigma,\rho)=2\int_{\mathbb{R}^{3}}\sigma V_{\rho}\,dx$
Consider $T_{J}(\rho)=\frac{J^{2}}{2I(\rho)}$, when $t$ is small enough, we have $\int_{\mathbb{R}^{3}}(\rho+t\sigma)\,dx>0$, thus $I(\rho+t\sigma)>0$. and $T_{J}(\rho+t\sigma)$ is well-defined. Easy to check the mass of $t\sigma$ is $m(t\sigma)=tm(\sigma)$, the center of mass is $\bar{x}(t\sigma)=\bar{x}(\sigma)$, and the moment of inertia is $I(t\sigma)=tI(\sigma)$. Thanks to Lemma 2.15, we can expand $I(\rho+t\sigma)$ and get
@@
$I(\sigma)+m(\sigma)\cdot r^{2}(\bar{x}(\rho)-\bar{x}(\sigma))=$ $\int_{\mathbb{R}^{3}}\sigma\cdot\left(r^{2}(x-\bar{x}(\sigma))+r^{2}(\bar{x}(\rho)-\bar{x}(\sigma))\right)\,dx$
$=$ $\int_{\mathbb{R}^{3}}\sigma\cdot(\langle x,x\rangle_{2}-2\langle\bar{x}(\sigma),x\rangle_{2}+\langle\bar{x}(\sigma),\bar{x}(\sigma)\rangle_{2}$
@@
$-2\langle\bar{x}(\rho),x\rangle_{2}+2\langle\bar{x}(\rho),x\rangle_{2})\,dx$
$=$ $\int_{\mathbb{R}^{3}}\sigma\cdot(r^{2}(x-\bar{x}(\rho))+2\langle\bar{x}(\sigma),\bar{x}(\sigma)\rangle_{2}$
$-2\langle\bar{x}(\sigma),x\rangle_{2}-2\langle\bar{x}(\sigma),\bar{x}(\rho)\rangle_{2}+2\langle\bar{x}(\rho),x\rangle_{2})\,dx$
$=$ $\int_{\mathbb{R}^{3}}\sigma\cdot r^{2}(x-\bar{x}(\rho))\,dx+2m(\sigma)\langle\bar{x}(\sigma),\bar{x}(\sigma)\rangle_{2}-2m(\sigma)\langle\bar{x}(\sigma),\bar{x}(\sigma)\rangle_{2}$
@@
$=$ $\int_{\mathbb{R}^{3}}r^{2}(x-\bar{x}(\rho))\cdot\sigma\,dx$
Therefore, $\frac{T_{J}(\rho+t\sigma)-T_{J}(\rho)}{t}\rightarrow\int_{\mathbb{R}^{3}}\frac{-J^{2}}{2I^{2}(\rho)}\cdot r^{2}(x-\bar{x}(\rho))\cdot\sigma\,dx$ as $t\rightarrow 0$. Collecting the results above, we finish the proof of Lemma 5.3. ∎
@@
The admissible perturbations are a bit different from $P_{0}(\rho)$ mentioned in [5, Section 4]. There the energy related to momentum is not $T_{J}(\rho)$ but $\frac{1}{2}\int_{\mathbb{R}^{3}}\rho(x)L\left(m_{\rho}(r(x))\right)r^{-2}(x)\,dx$, in which the term $r^{-2}(x)$ will be singular when $r(x)\rightarrow 0$. Therefore, the authors require $\sigma(x)=0$ when $r(x)$ is small.
We are now ready to prove the differentiability of $E_{J}$ at local minimizers:
@@
If $\rho$ minimizes $E_{J}(\rho)$ locally on ${R}\left(\mathbb{R}^{3}\right)$, then $E_{J}(\rho)$ is finite, $\rho\in W$. In particular, by Lemma 5.3, we know $E_{J}(\rho)$ is $P_{\infty}(\rho)$-differentiable at $\rho$, and the derivative at $\rho$ is $E_{J}^{\prime}(\rho)$ given in (29).
@@
Since $\int_{\mathbb{R}^{3}}\rho\,dx=1>0$, similar to Remark 5.1 we know $T_{J}(\rho)$ and $G(\rho,\rho)$ are finite. Since $\rho$ is a local minimizer, $\exists\delta>0$, such that $\forall\sigma\in{R}\left(\mathbb{R}^{3}\right)$ with $W^{\infty}(\rho,\sigma)<\delta$, one has $E_{J}(\rho)\leq E_{J}(\sigma)$. Thanks to Lemma 4.5 (vivi), take $\delta$ small enough, we can find a $\sigma\in{R}\left(\mathbb{R}^{3}\right)\cap L^{\infty}\left(\mathbb{R}^{3}\right)$, with $W^{\infty}(\rho,\sigma)<\delta$ and $\int_{\mathbb{R}^{3}}\sigma\,dx>0$. Similar as above we have $G(\sigma,\sigma)<\infty,T_{J}(\sigma)<\infty$. Thanks to Lemma 5.2, we know $U(\sigma)<\infty$, and then $E_{J}(\rho)\leq E_{J}(\sigma)<\infty$, which implies $U(\rho)<\infty$. Then we have $\rho\in W$. ∎